# DROID: A Large-Scale In-The-Wild Robot Manipulation Dataset https://droid-dataset.github.io

### v1

@@ Abstract
††footnotetext: Project Co-leads, correspondence to alexkhaz@stanford.edu, pertsch@berkeley.eduThe creation of large, diverse, high-quality robot manipulation datasets is an important stepping stone on the path toward more capable and robust robotic manipulation policies. However, creating such datasets is challenging: collecting robot manipulation data in diverse environments poses logistical and safety challenges and requires substantial investments in hardware and human labour. As a result, even the most general robot manipulation policies today are mostly trained on data collected in a small number of environments with limited scene and task diversity. In this work, we introduce DROID (Distributed Robot Interaction Dataset), a diverse robot manipulation dataset with 76k demonstration trajectories or 350 hours of interaction data, collected across 564 scenes and 86 tasks by 50 data collectors in North America, Asia, and Europe over the course of 12 months. We demonstrate that training with DROID leads to policies with higher performance and improved generalization ability. We open source the full dataset, policy learning code, and a detailed guide for reproducing our robot hardware setup.
@@ I Introduction
††footnotetext: Affiliations: Stanford University; University of California, Berkeley; Toyota Research Institute; Carnegie Mellon University; University of Texas, Austin; University of Montreal; University of Edinburgh; Princeton University; University of Washington; Korea Advanced Institute of Science & Technology (KAIST); University of California, San Diego; Google DeepMind; University of California, Davis; University of Pennsylvania; Columbia University| Dataset | # Traj. | # Verbs | # Scenes | Lang. Instruct. | Cam. Calibration | Public Robot | Collection |
@@
| DobbE [39] | 5.6k | 6 | 216 | ✓ | n/a | (✓) | human tool-based |
| Open X-Embodiment [33] | 1.4M | 217 | 311 | (✓) | ✗ | (✓) | dataset aggregation |
@@ ACKNOWLEDGMENT
We thank the Toyota Research Institute (TRI) for their support in various aspects of this project, from data collection to compute for policy training. This work was supported by the Google TPU Research Cloud. We further acknowledge the following funding sources: Chelsea Finn’s group was supported by TRI and ONR grants N00014-20-1-2675 and N00014-22-1-2621; Sergey Levine’s group was supported by TRI, NSF FRR IIS-2150826, and ONR N00014-20-1-2383; Ram Ramamoorthy’s group was supported by the United Kingdom Research and Innovation through grant EP/S023208/1 to the EPSRC Centre for Doctoral Training in Robotics and Autonomous Systems (RAS) and grant EP/V026607/1 to the UKRI Research Node on Trustworthy Autonomous Systems Governance and Regulation; Dorsa Sadigh’s group was supported by TRI and ONR grant N00014-22-1-2293; Glen Berseth’s group acknowledges funding support from NSERC and CIFAR and compute support from Digital Research Alliance of Canada, Mila IDT and NVidia; Jeannette Bohg’s group was supported by TRI, Intrinsic, Toshiba and the National Science Foundation under Grant 2327974.

### v2

@@ Abstract
††footnotetext: ∗Project Co-leads, correspondence to alexkhaz@stanford.edu, pertsch@berkeley.eduThe creation of large, diverse, high-quality robot manipulation datasets is an important stepping stone on the path toward more capable and robust robotic manipulation policies. However, creating such datasets is challenging: collecting robot manipulation data in diverse environments poses logistical and safety challenges and requires substantial investments in hardware and human labour. As a result, even the most general robot manipulation policies today are mostly trained on data collected in a small number of environments with limited scene and task diversity. In this work, we introduce DROID (Distributed Robot Interaction Dataset), a diverse robot manipulation dataset with 76k demonstration trajectories or 350 hours of interaction data, collected across 564 scenes and 86 tasks by 50 data collectors in North America, Asia, and Europe over the course of 12 months. We demonstrate that training with DROID leads to policies with higher performance and improved generalization ability. We open source the full dataset, policy learning code, and a detailed guide for reproducing our robot hardware setup.
@@ I Introduction
††footnotetext: Affiliations: 1Stanford University; 2University of California, Berkeley; 3Toyota Research Institute; 4Carnegie Mellon University; 5University of Texas, Austin; 6University of Montreal; 7University of Edinburgh; 8Princeton University; 9University of Washington; 10Korea Advanced Institute of Science & Technology (KAIST); 11University of California, San Diego; 12Google DeepMind; 13University of California, Davis; 14University of Pennsylvania; 15Columbia University; 16Yonsei University| Dataset | # Traj. | # Verbs | # Scenes | Lang. Instruct. | Cam. Calibration | Public Robot | Collection |
@@
| DobbE [47]∗ | 5.6k | 6 | 216 | ✓ | n/a | (✓) | human tool-based |
| Open X-Embodiment [39]† | 1.4M | 217 | 311 | (✓) | ✗ | (✓) | dataset aggregation |
@@ III-B Data Collection Protocol
Since the initial extrinsic calibration parameters, provided through conventional calibration detailed above, may not always be accurate due to factors such as checkerboard misalignment, inconsistent lighting, or errors inherent to the OpenCV calibration method, we address these inaccuracies in Appendix G. We discuss in detail the automatic post-hoc calibration process and provide three comprehensive sets of camera calibration matrices for the DROID dataset, each accompanied by respective quality assessment metrics. These include camera-to-base calibrations for around 36k unique scenes with one camera calibrated relative to the base, camera-to-camera calibrations for all scenes, and a curated superset of 24k scenes covering all three calibration methods with both cameras calibrated relative to the base. These refined calibrations enhance the dataset’s suitability for robust geometric understanding in robotics and 3D perception tasks. For more details, please see Sec. Appendix G.
@@ ACKNOWLEDGMENT
We thank the Toyota Research Institute (TRI) for their support in various aspects of this project, from data collection to compute for policy training. This work was supported by the Google TPU Research Cloud. We further acknowledge the following funding sources: Chelsea Finn’s group was supported by TRI and ONR grants N00014-20-1-2675 and N00014-22-1-2621; Sergey Levine’s group was supported by TRI, NSF FRR IIS-2150826, and ONR N00014-20-1-2383; Ram Ramamoorthy’s group was supported by the United Kingdom Research and Innovation through grant EP/S023208/1 to the EPSRC Centre for Doctoral Training in Robotics and Autonomous Systems (RAS) and grant EP/V026607/1 to the UKRI Research Node on Trustworthy Autonomous Systems Governance and Regulation; Dorsa Sadigh’s group was supported by TRI and ONR grant N00014-22-1-2293; Glen Berseth’s group acknowledges funding support from NSERC and CIFAR and compute support from Digital Research Alliance of Canada, Mila IDT and NVidia; Jeannette Bohg’s group was supported by TRI, Intrinsic, Toshiba and the National Science Foundation under Grant 2327974; Joseph Lim’s group was supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grants (No.2019-0-00075, Artificial Intelligence Graduate School Program, KAIST; No.2022-0-00077, AI Technology Development for Commonsense Extraction, Reasoning, and Inference from Heterogeneous Data), and a National Research Foundation of Korea (NRF) grant (NRF-2021H1D3A2A03103683) funded by the Korean government (MSIT).
@@ F-B Training Batch Construction
Appendix G Automatic Camera Calibration
In this section, we provide three comprehensive sets of camera calibration matrices for the DROID dataset with their respective quality assessment metrics, including camera-to-base calibrations for 36k unique scenes with one of the cameras calibrated with respect to base, camera-to-camera calibrations for all scenes, and a curated superset of 24k scenes encompassing all three methods and with both cameras calibrated with respect to base, facilitating downstream robust geometric understanding in robotics and 3D perception tasks.
Accurate camera calibration can be very useful in robotics and 3D perception, as it enables the consistent encoding of spatial geometry from visual data. It serves as the backbone for various downstream tasks in robotics manipulation, such as learning viewpoint invariant representations [6, 56] or grounding actions through 3D vision-and-language models [52, 66], thus enabling the robotics agents to achieve geometric and visual generalization. In robotic applications, calibration allows for precise scene understanding and interaction by aligning sensor observations to a shared spatial frame.
The DROID dataset provides initial extrinsic parameters (Sec. III) that transform coordinates from the camera frame to the robot base frame. However, these calibrations are not always accurate, primarily due to slight errors that can arise during the manual calibration process, such as imperfect checkerboard placements, variations in lighting conditions, or inaccuracies in the OpenCV calibration procedure performed at the start of each data collection session. Following the data collection efforts outlined in Sec. III, we additionally focus on providing robust calibration values for the collected dataset in an off-line post-hoc manner. This process utilizes recent advances in deep-learning based perception systems [30, 61, 41, 34] to automatically calibrate the relevant cameras and provide quality metrics in a post-hoc manner.
The following sections details the automatic post-hoc calibration of the pre-collected DROID dataset. It focuses on two key types of calibration: (i) camera-to-robot base extrinsic calibration, which computes the transformation between a fixed camera and the robot’s kinematic base; and (ii) camera-to-camera extrinsic calibration, which estimates the relative pose i.e. orientation and translation between two external cameras. Both are essential for fusing multi-view observations as well as allowing the grounding of robotics actions in 3D, thus enabling spatially grounded robotic behaviors. This section is divided into 3 sub-sections. We first detail the quality metric assessment of existing camera-to-robot base calibration (Sec. G-A) provided after the data-collect phase in Sec. III. This would allow us to filter the extrinsics provided during data-collect and provide certain guarantees regarding the calibration already provided. We then explain how to calibrate additional cameras with respect to the base using a fully automatic pipeline [34] while also providing guarantees in terms of quality metric i.e. reprojection error (Sec. G-B). Furthermore, we discuss calibrating external cameras with respect to each other in Sec. G-C. Finally, we include a discussion on limitations and future work in Sec. G-D.
G-A Quality Assessment of Existing Camera-to-Robot Base Calibration
To evaluate the quality of the existing camera-to-robot base transformation (), we project known 3D keypoints , obtained via forward kinematics for given joint angles , into the image plane using the extrinsic matrix and camera intrinsics . The 2D projections are computed as , where denotes perspective projection followed by normalization.
These 2D keypoints are used to guide a Segment-Anything (SAM) [29] instance segmentation model, which predicts masks . Simultaneously, synthetic robot masks are rendered using PyTorch3D by importing the robot’s mesh geometry and kinematic structure defined in its URDF. Each joint angle configuration is applied to the URDF to compute the articulated 3D mesh pose of the robot. The resulting mesh is transformed to the camera frame using the same extrinsic transformation . The posed mesh is then rasterized into a binary silhouette using a differentiable renderer with the corresponding camera intrinsics . This rendered mask serves as the ground-truth projection for evaluating the alignment quality of the predicted segmentation.
We compute the Intersection-over-Union (IoU) between the predicted and ground-truth masks as . Only SAM masks with confidence scores greater than 0.65 are retained. A final threshold of is used to identify high-quality projections, filtering out poorly aligned frames. We report the mean IoU across 5 equally subsampled frames in a video sequence as a measure of calibration quality. Using this process, we identified a total of around 30k scenes with either the left or right camera well calibrated with respect to the scene. The whole process took around 1 day on 8-A100 Nvidia-GPUs.
G-B Automatic Camera-to-Robot Base Calibration
To supplement the filtering strategy outlined in Sec. G-A and bring in additional cameras for the camera-to-robot base calibration, we additionally ran a tuned version of CtRNet-X [34] out-of-the-box on all of DROID dataset. We used the original codebase provided by the authors as well as the hyperparameters tuned for the DROID dataset. CtRNet-X is a feed-forward approach that detects keypoints on robot using a neural-network and matches it with ground-truth 3D keypoint trajectory in the video. Additionally, they utilize a CLIP [41]-guided robot part detection to dynamically select visible keypoints. Following author’s implementation, we use a confidence threshold of 0.08, CLIP [41] models’ end-effector confidence of 0.1 and robot base confidence of 0.05. To evaluate the quality of our camera-to-base calibration, we compute the reprojection error between detected 2D keypoints and their corresponding 3D projections using the estimated camera pose and intrinsics. To ensure robustness, we first discard low-confidence 2D observations based on a fixed threshold. We then apply a Median Absolute Deviation (MAD) based outlier rejection strategy: we calculate the median of all reprojection errors, compute the absolute deviation of each error from the median, and identify inliers as those within 2.5 times the MAD. This robust statistical filtering helps suppress the influence of large outliers, leading to a more reliable estimate of the mean reprojection error. For the final filtering, we select scenes with a mean reprojection-error of 20 or less.
Through this process, we identified a total of around 12k scenes which have either the left or right camera correctly calibrated with respect to the base. This process took around 5 days on 8-A100 Nvidia GPUs. Since the number of well-calibrated scenes overlapped between the two strategies, we are able to calibrate around 36k unique scenes with either the left or right camera well calibrated with respect to the scene using the strategy outlined in Sec. G-A and Sec. G. Figure 11 shows randomly selected scenes with synthetically rendered robot masks using PyTorch3D [44], qualitatively demonstrating the high accuracy of our filtered camera-to-robot base calibration from both the above-mentioned approaches. Furthermore, we present the distribution of IoU and reprojection errors after applying the improved calibration strategy and filtering. These results are shown in Fig. 15.
G-C Automatic Camera-to-Camera Calibration
We utilize the recently released DUSt3R [61] framework for improved Camera-to-Camera calibration. DUSt3R [61] supports both relative and absolute pose estimation. For relative pose estimation, DUST3R proposes obtaining 2D–3D correspondences between a query image and a reference image , followed by PnP-RANSAC [15, 31] using known or estimated intrinsics. The relative pose between and can also be converted to an absolute pose in world coordinates by aligning predicted pointmaps to a known scale, typically via a ground truth pointmap for . However, this approach still requires scale alignment post-optimization and can suffer from ambiguities due to noise or uncertainty in the predicted geometry.
We modify the pose optimization pipeline of DUSt3R [61] to utilize depth maps and known camera intrinsics as an input in the optimization pipeline to recover absolute poses in a consistent metric scale. Specifically, we begin by running DUSt3R inference on image pairs to extract dense 3D pointmaps. These predicted pointmaps are aligned to ground truth 3D point clouds—constructed from depth and intrinsics—to compute a global scale factor. We then perform a global optimization step, where the ground truth depth and intrinsics are fixed, and camera poses are refined to minimize the 3D alignment error across the scene. This approach enables accurate, scale-aware absolute pose estimation without relying on post-hoc scale alignment.
By fixing the depth and intrinsics during optimization, we ensure that the recovered poses are globally consistent and metrically accurate. Importantly, our method operates directly on unmodified DUSt3R [61] outputs, requiring no additional training or manual scale correction. Figures 12 and 13 show qualitative improvements in point cloud alignment following this optimization step.
To assess the quality of the recovered camera-to-camera calibration, we report the number of matched points between views, following the original implementation by [61]. For each image pair, we extract high-confidence 3D points, project them into 2D using known intrinsics, and identify reciprocal nearest neighbors in 3D space as reliable matches. Formally, given pointmaps and , we define the match set . In practice, we qualitatively observe that a higher number of reciprocal 3D matches visually correlates with the geometric quality of estimated poses (see Fig. 12). Although the number of matches serves as a reasonable proxy for assessing the quality of estimated poses, we observed some false positives in visually cluttered scenes. To enhance robustness, one could lower the filtering threshold or incorporate the quality assessment described in Sec. G-A to further refine the filtering process. Figure 14 shows the distribution of match counts across labs. While some labs exhibit strong geometric consistency, others struggle due to challenging conditions like clutter or poor lighting. For all videos, the first frame was used for pose refinement via our modified DUSt3R [61] pipeline. Future improvements could include ensembling predictions across frames or leveraging temporal consistency to further stabilize pose estimation or finetuning DUST3R-like methods on table-top cluttered datasets observed in robotics manipulation settings.
G-D Limitations and Future Work
Calibrating a large-scale dataset like DROID is a challenging task. To ensure accuracy and provide guarantees at each step, we divided the calibration process into three distinct stages. Since the complete process is fully automatic, there are still some false positives and future work could look at further improving on these inconsistencies. Part of our camera-to-base calibration relied on running an out-of-the-box model which was trained on Franka panda robot. While successful, its zero-shot generalizability to other robots (without requiring further training or finetuning) remains to be seen.
Future work could look at using foundation models to segment out the robot or gripper and estimating keypoints on specific parts of the robot. This could provide a more generalizable solution that could be readily applied to any robot collected data in-the-wild. Our Camera-to-Camera calibration relied on a recent paradigm in 3D deep learning, namely the prediction of point-maps. Despite using a modified version of DUSt3R [61], which utilized privileged depth information for pose optimization, it relied on the original checkpoints provided by the authors and hence also borrowed the limitations of the original model. Despite decent success, the model at times fails on scenes with clutter and challenging table-top settings with little to no overlap between images. As the quality of these models keeps improving [62, 60], we believe it would be a valuable direction to leverage these improved models for more robust camera-to-camera calibration, particularly in cluttered and low-overlap scenarios where traditional feature matching or earlier models struggle.
G-E Conclusion
The approaches outlined in each of the aforementioned sections i.e. Sec. G-A, Sec. G-B and Sec. G-C have different guarantees in terms of quality metrics. Hence, for the final calibration release we offer 3 different sets of camera calibration matrices. The first one contains Camera-to-Robot base calibration from 36k unique scenes with either the left or right camera calibrated with respect to the robot base. This includes results after the combined calibration methods outlined in Sec. G-A and Sec. G-B. The second set of calibration includes all Camera-to-Camera calibration matrices, i.e. the relative transformations for all scenes in DROID dataset using the approach outlined in Sec. G-C. Finally, we also release a third set of calibration with includes a superset of all methods which, totaling around 24k scenes with both cameras calibrated with respect to the base and with a mix of 3 different approaches and individual guarantees. In this superset, we use an IOU threshold of 0.6, a reprojection error of 20 and top 30% quantile based on number of matches for each stage of calibration described earlier in Sections G-A, G-B and G-C respectively. We hope this effort is useful for 3D vision and robotics manipulation research and also serves an inspiration for off-line automatic camera calibration of in-the-wild robotics manipulation datasets.